OpenStudy (anonymous):

GRAPH -1/3x -1/4y=1/2 steps plz

OpenStudy (aum):

To get rid of the fractions, multiply throughout by 12 which is the least common denominator of 3, 4 and 2. What do you get?

OpenStudy (aum):

-1/3x - 1/4y = 1/2 multiply by 12 to get rid of fractions: -4x - 3y = 6 multiply by -1 to get rid of negative sign on the left: 4x + 3y = -6 This is a straight line. To graph it all you need is 2 points. Put x = 0 in the equation and find y. That will be the y-intercept. Mark this point on the graph on the y-axis. Then put y = 0 and find x. That will be the x-intercept. Mark this point on the graph on the x-axis. Draw a straight line passing through the two points.

OpenStudy (anonymous):

Thank you so much I couldn't find a solution as to how I would attempt to do this problem with fractions

OpenStudy (aum):

You are welcome.